OpenStudy (anonymous):

A company produces baseballs at 3 different plants: plant A, plant B and plant C. doubling plant A's production is equal to 1/3 of the company's total production. doubling plant B's production is equal to the company's total production.tripling plant C's production is equal to the company's total production what us the ratio of plant A's production to plant B's production to plant C's production?

OpenStudy (anonymous):

okay here we go. this is just a bunch of logic, really. so when plant A is doubled, its equal to 1/3 of plant b. and when plant B is doubled, that number is equal to the total output of the company. so... 2a=1/3 a=1/6 of total production. 2b= 1 b=1/2=3/6 total production. 3c=1 c=1/3=2/6 so a/b/c= 1 to 3 to 2

OpenStudy (anonymous):

thank u so much...yep its question of SAT...so its so logical...

OpenStudy (anonymous):

hahaha no problem :D i hated the SATs good luck!